# Energy-Efficient Channel Decoding for Wireless Federated Learning: Convergence Analysis and Adaptive Design

### v2

@@ Abstract
One of the most critical challenges for deploying distributed learning solutions, such as federated learning (FL), in wireless networks is the limited battery capacity of mobile clients. While it is a common belief that the major energy consumption of mobile clients comes from the uplink data transmission, this paper presents a novel finding, namely the channel decoding operation also contributes significantly to the overall energy consumption of mobile clients in FL. Motivated by this new observation, we propose an energy-efficient adaptive channel decoding scheme that leverages the intrinsic robustness of FL to model errors. In particular, the robustness is exploited to reduce the energy consumption of channel decoders at mobile clients by adaptively adjusting the number of decoding iterations. We theoretically prove that wireless FL with communication errors can converge at the same rate as the case with error-free communication as long as the bit error rate (BER) is properly constrained. An adaptive channel decoding scheme is then proposed to improve the energy efficiency of wireless FL systems. Experimental results demonstrate that the proposed method maintains the same learning accuracy while reducing the channel decoding energy consumption by 20% when compared to existing approaches.
@@ I-A Motivation
Deep learning has achieved remarkable success in numerous realms, such as computer vision and natural language processing [1]. However, the conventional centralized learning paradigm encounters the data isolation issue due to the rising concern about privacy [2]. As a result, uploading the distributed data, which may contain privacy-sensitive information, for centralized learning becomes prohibited. To address this issue, federated learning (FL) was proposed [3] to enable the distributed training of deep learning models without sharing the data of users (e.g., IoT devices). However, to fully unlock the potential of FL, we need to tackle several critical challenges, such as data and device heterogeneity [4, 5, 6], as well as the communication bottleneck[7]. Another issue of wireless FL is the energy consumption of mobile clients due to heavy computation and communication workloads for training and transferring the deep neural network (DNN) parameters[8].
@@
Many innovative approaches have been proposed to reduce the energy consumption of mobile clients in FL from different perspectives. For instance, reinforcement learning was utilized in [9] to identify near-optimal client selection and resource allocation to reduce both the convergence time and energy consumption of mobile clients. In [10], the trade-off between learning accuracy and total energy consumption of users was optimized through computational and communication resource allocation. In [11], the transmission and computation energy was minimized by bandwidth allocation and power control. Besides system-level resource management, techniques such as model sparsification [12] and quantization [13, 14, 15, 16, 17, 18] have been proposed to minimize the communication overhead, thus reducing the energy consumption of the transmitters. Authors of [19] minimized the transmission energy consumption of mobile clients while maintaining the learning performance by optimizing the transmit power, user selection and uplink resource allocation. Authors of [20] adopted quantization and periodic aggregation to reduce the transmission volume and energy, which also provided a theoretical framework about the impact of quantization and periodic aggregation on the convergence rate of FL. Model sparsification and quantization techniques are also beneficial for minimizing the processing energy consumption [21, 8]. Authors of [21] proposed to introduce sparsity at training time to reduce the on-device training energy. Authors of [8] reduced the local computation energy by quantization on both DNN weights and activations.
@@
However, most works on energy-efficient FL focused on the energy consumption of mobile transmitters while few of them paid attention to the energy consumption of channel decoders in mobile receivers. As will be shown in Section II-B, channel decoders consume a substantial proportion of energy because of the iterative processes to decode massive DNN parameters in each communication round of wireless FL.
@@
To address the research gap, in this work, we first quantitatively analyze the energy consumption of different components in a mobile unit and show that the channel decoder contributes significantly to the overall energy consumption. Building on this observation, an energy-efficient adaptive channel decoding scheme is proposed for wireless FL systems, by leveraging the fact that FL is robust to model errors due to its statistical learning nature[22]. Furthermore, by extending the convergence analysis in[20] to the scenario of FL with downlink communication erros, we investigate the relationship between the bit error rate (BER) and the convergence rate of wireless FL, and determine the maximum tolerable BER that can maintain the same convergence rate as the case with error-free communication. Finally, we propose a practical approach to adjust the number of channel decoding iterations in each communication round to achieve the maximum tolerable BER, while reducing the energy consumption of mobile clients in wireless FL.
@@ I-B Contributions
  • We theoretically analyze the relationship between BER and the convergence rate of wireless FL. Specially, an upper bound of the average expected gradient norm of the global training loss with respect to BERs over communication rounds is derived by extending the analysis in [20], based on which, the maximum tolerable BER to achieve the same convergence rate as error-free FL is determined.
  • Based on the theoretical analysis, an adaptive low-density parity checking (LDPC) decoding mechanism is proposed, which controls the maximum number of decoding iterations to satisfy the BER requirement in every communication round and maintain the target convergence rate as error-free wireless FL, while reducing the energy consumption of channel decoders in wireless FL.
  • Experimental results on both IID and non-IID datasets show that the proposed adaptive scheme effectively reduces the energy consumption of channel decoders by 20% when compared with the traditional schemes which set a fixed maximum number of channel decoding iterations, while guaranteeing the same learning performance.
@@ I-C Organization
The remainder of this paper is organized as follows. Section II introduces the system model and presents a comprehensive energy consumption analysis for mobile units. Section III develops an energy-efficient adaptive LDPC decoding scheme for wireless FL systems by exploiting the relationship between the convergence rate and the achieved BER of LDPC decoding. Experimental results are presented in Section IV. Finally, conclusions are drawn in Section V.
II System Model and Preliminaries
@@
$\mathbf{w}$ DNN model
@@
$f$ The loss function of training
$f^{*}$ Minimum value of the global training loss
$R$ The total communication rounds
@@
$\sigma_{L}^{2}$ Variance of local gradient estimator
$\sigma_{G}^{2}$ Dissimilarity between local and global gradient
$n$ Number of total clients
$N$ Each DNN parameter is with $N$ bit representation
$d$ Dimension of the DNN model
$T$ Total training iterations, i.e., $T=RE$.
$L$ The $L$-smooth parameter
@@
$Q$ Maximum number of LDPC decoding iterations
$M_{r}$ Range of the global model parameters in the $r$-th communication round
$\hat{M}$ Maximum range of global model parameters for all communication rounds
(a) Wireless FL system: Multiple mobile clients are coordinated by one remote server through noisy wireless channels, where model updates and model weights are communicated in the uplink and downlink, respectively.
(b) System operations: The server performs model aggregation, digitalization, LDPC encoding, and model broadcasting. Mobile clients perform LDPC decoding, local training, and uploading the model updates.
Figure 1: Wireless FL in digital communication network and its operations.
@@ II-A Wireless FL
Fig. 1(a) depicts a wireless FL system, where model training of different clients is coordinated by a remote server. The FL problem can be formulated as follows [23, 24, 25]:
$\min_{\mathbf{w}\in{\mathbb{R}^{d}}}f(\mathbf{w})\triangleq\frac{1}{n}\sum_{k=1}^{n}f_{k}(\mathbf{w}),$ (1)
@@
where $\mathbf{w}\in{\mathbb{R}^{d}}$ denotes the DNN model with $d$ parameters, $n$ represents the number of mobile clients, $f_{k}\left(\cdot\right)$ is the local loss function of the $k$-th client, and $f\left(\cdot\right)$ is the global loss function.
Figure 2: Digitalization scheme for model parameters. This part gives an example, say [min, max] is digitalized by 3-bits and shows how each value inside is mapped to ”000”, ”001”, ”…”, ”111”.
@@
To train a shared global model, the server communicates with the clients for multiple rounds through noisy wireless channels. As shown in Fig. 1(b), in each communication round, the global model at the server is broadcast to the clients. Each client then performs local training based on the received global model and transmits the model updates after local training to the server for aggregation. Specifically, in the $r$-th communication round, the global model is digitalized by $N$ bits, encoded by a channel code, and broadcast to clients. The model parameter digitalization scheme is illustrated in Fig. 2, where the range of [min, max] is divided into $12^{N}-$ 1 intervals and the width of each interval is $\Delta=\frac{\text{max-min}}{2^{N}-1}$. In this paper, we assume $N$ is large enough so the continuous and digital representations of model parameters have negligible differences. The model parameters received by the $k$-th client in the $r$-th communication round is given as follows:
$\mathbf{w}^{k}_{r,0}=\mathbf{\tilde{w}}_{r},\ k=1,\cdots,n,\ r=0,\cdots,R-1,$ (2)
@@
where $\mathbf{\tilde{w}}_{r}$ denotes the distorted version of $\mathbf{w}_{r}$ due to channel decoding errors. The impact of channel decoding errors can be quantified by considering the different error bit positions in the $N$-bit representations of model parameters, which will be analyzed in Section III-A. Each client updates the received global model by conducting $E$ steps of local mini-batch Stochastic Gradient Descent (SGD) with a learning rate $\eta$ as follows:
$\begin{split}\mathbf{w}^{k}_{r,e+1}&=\mathbf{w}^{k}_{r,e}-\eta{\nabla}f_{k}(\mathbf{w}^{k}_{r,e};\xi_{r,e}^{k}),\\ &e=0,\cdots,E-1,\ k=1,\cdots,n,\end{split}$ (3)
@@
where ${\nabla}f_{k}(\mathbf{w}^{k}_{r,e};\xi_{r,e}^{k})$ denotes the gradient computed from a mini-batch datasets $\xi_{r,e}^{k}$ at the $k$-th client.
By completing $E$ steps of local training, the local model update of the $k$-th client can be expressed as follows:
$\begin{split}\Delta\mathbf{w}^{k}_{r}&=\mathbf{w}^{k}_{r,E}-\mathbf{w}^{k}_{r,0},\\ &r=0,\cdots,R-1,\ k=1,\cdots,n,\end{split}$ (4)
@@
$\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{n}\sum_{k=1}^{n}\Delta\mathbf{\tilde{w}}^{k}_{r},\ r=0,\cdots,R-1.$ (5)
In this work, we focus on the downlink communication rather than the uplink since our objective is to design novel channel decoding schemes to save the energy consumption of mobile clients. For the uplink, the channel decoding happens at the server, where a powerful channel decoder as well as a sufficient number of decoding iterations can always be applied to achieve satisfactory BER performance. As such, we consider $\Delta\mathbf{\tilde{w}}_{r}^{k}=\Delta\mathbf{w}_{r}^{k}$ in the following.
@@ II-B Energy Consumption Evaluation
There are two main energy-consuming operations for clients participating in wireless FL, namely model training and model communication. For the latter, it is believed that the energy is mostly consumed by the data transmission. In this section, we evaluate the energy consumption of different components, taking the training of AlexNet for image classification tasks as an example [26]. The analysis is based on actual measurements obtained from fabricated chips published in recent years [27, 28, 29]. For fair comparisons, all data are collected for chips fabricated in the 65 nm complementary metal-oxide semiconductor (CMOS) process.
In[27], a fabricated AI hardware accelerator was reported to achieve the energy efficiency of 13.7 mJ/epoch for 8-bit quantized AlexNet. With five epochs of local training, the processing energy consumption in each communication round is 68.5 mJ. In [28], a fifth-generation (5G) New Radio wireless transmitter and receiver were reported to achieve the energy efficiencies of 47.9 pJ/b and 33.3 pJ/b for transferring uncoded data, respectively, under a communication distance of around 200 meters. Therefore, the energy required to transmit and receive the 8-bit AlexNet with 60 million parameters in a single communication round is about 39 mJ. Assuming a code rate of 1/2, the communication energy consumption for coded transmission of the AlexNet model parameters in each round is 78 mJ. Besides, the LDPC decoder in [29] achieves the energy efficiency of 20.1 pJ/b/iteration, which was measured with 10 decoding iterations on average. Accordingly, the energy consumption of the LDPC decoder for decoding the AlexNet model parameters is 96.5 mJ. The detailed comparison is presented in Table II, which indicates that the energy consumption of LDPC decoding accounts for a large proportion of the overall energy consumption at clients in wireless FL.
TABLE II: Energy Consumption of Main Components at a Mobile Unit in Wireless FL for Training an AlexNet
@@
RF Transceiver 81.2 pJ/b[28] 78 mJ
LDPC Decoder 20.1 pJ/b/iter[29] 96.5 mJ

### v3

@@ Abstract
One of the most critical challenges for deploying distributed learning solutions, such as federated learning (FL), in wireless networks is the limited battery capacity of mobile clients. While it is a common belief that the major energy consumption of mobile clients comes from the uplink data transmission, this paper presents a novel finding, namely channel decoding also contributes significantly to the overall energy consumption of mobile clients in FL. Motivated by this new observation, we propose an energy-efficient adaptive channel decoding scheme that leverages the intrinsic robustness of FL to model errors. In particular, the robustness is exploited to reduce the energy consumption of channel decoders at mobile clients by adaptively adjusting the number of decoding iterations. We theoretically prove that wireless FL with communication errors can converge at the same rate as the case with error-free communication provided the bit error rate (BER) is properly constrained. An adaptive channel decoding scheme is then proposed to improve the energy efficiency of wireless FL systems. Experimental results demonstrate that the proposed method maintains the same learning accuracy while reducing the channel decoding energy consumption by $\sim$ 20% when compared to an existing approach.
@@ I-A Motivation
Deep learning has achieved remarkable successes in numerous realms, such as computer vision and natural language processing [2]. However, the conventional centralized learning paradigm encounters the data isolation issue due to the rising concern about privacy [3]. As a result, uploading the distributed data, which may contain privacy-sensitive information, for centralized learning becomes prohibited. To address this issue, federated learning (FL) was proposed [4] to enable the distributed training without sharing the data of users (e.g., IoT devices). However, to fully unlock the potential of FL, we need to tackle several critical challenges, such as data and device heterogeneity [5, 6], as well as the communication bottleneck[7]. Another issue of wireless FL is the energy consumption of mobile clients due to heavy computation and communication workloads for training deep neural networks (DNNs) and transferring model parameters[8].
@@
Many innovative approaches have been proposed to reduce the energy consumption of mobile clients in wireless FL from different perspectives. For instance, reinforcement learning was utilized in [9] to determine the near-optimal client selection and resource allocation scheme to reduce both the convergence time and energy consumption of mobile clients. In [10], the trade-off between learning accuracy and total energy consumption of clients was optimized through computational and communication resource allocation. In [11], the transmission and computation energy was minimized under a latency constraint by optimizing the time, bandwidth allocation, power control, computation frequencies, and learning accuracy. Also, authors of [12] minimized the transmission energy consumption of mobile clients while maintaining the learning performance by further designing the client selection strategies.
@@
Besides system-level resource management, techniques such as model sparsification [13] and quantization [14, 15] have been proposed to minimize the communication overhead, thus reducing the energy consumption of clients’ transmitters. Both quantization and periodic aggregation were adopted in [16] to reduce the communication overhead and energy consumption, where a theoretical framework about the impact of parameter quantization on the convergence rate was also provided. Model sparsification and quantization techniques are also beneficial for minimizing the DNN processing energy consumption at mobile clients of wireless FL [17, 18]. However, most works on energy-efficient FL focused on the energy consumption of mobile processors and transmitters, while few of them paid attention to the energy consumption of channel decoders in mobile receivers. As will be shown in Section II-B, channel decoders may consume a substantial proportion of energy because of the iterative processes to decode massive DNN parameters in each communication round of wireless FL.
@@
To address the research gap, in this work, we first quantitatively analyze the energy consumption of different components in a mobile unit and show that the channel decoder contributes significantly to the overall energy consumption. Building on this observation, an energy-efficient adaptive channel decoding scheme is proposed for wireless FL systems, by leveraging the fact that FL is robust to model errors due to its statistical learning nature[19]. Furthermore, by extending the convergence analysis in[16] to scenarios with downlink communication erros, we formalize the relationship between the bit error rate (BER) and the convergence performance of wireless FL. A decreasing rule of BER with respect to communications rounds is determined, which can maintain the same convergence rate as the case with error-free communication. Finally, we propose a practical approach to adjust the number of channel decoding iterations in each communication round to achieve the desired BER, while reducing the energy consumption of mobile clients in wireless FL.
@@ I-B Contributions
  • We theoretically analyze the relationship between BER and the convergence performance of wireless FL. Specially, an upper bound of the average expected gradient norm of the global training loss function with respect to BERs over communication rounds is derived by extending the analysis in [16]. Based on the analysis, a decreasing rule of BERs is proposed to achieve the same convergence rate as FL with error-free communication.
  • With the decreasing rule of BERs, an adaptive low-density parity checking (LDPC) decoding mechanism is proposed, which controls the maximum number of decoding iterations to achieve the desired BER requirement in every communication round. This provides a practical approach to maintain the target convergence rate, while reducing the channel decoding energy consumption of mobile clients in wireless FL.
  • Experimental results on both IID and non-IID datasets show that with a given number of communication rounds, the proposed adaptive LDPC decoding scheme effectively reduces the energy consumption of channel decoders by 20% when compared with a baseline scheme adopting a fixed maximum number of channel decoding iterations, while enjoying the same learning performance.
@@ I-C Organization
The remainder of this paper is organized as follows. Section II introduces the system model and presents a comprehensive energy consumption analysis for mobile clients. Section III develops an energy-efficient adaptive LDPC decoding scheme for wireless FL systems by exploiting the relationship between the convergence rate and the achieved BER of LDPC decoding. Experimental results are presented in Section IV. Finally, conclusions are drawn in Section V.
II System Model and Energy Consumption Evaluation
@@
$\mathbf{w}_{r}$ Global model in the $r$-th communication round
@@
$f\left(\cdot\right)$ The global loss function
$f^{*}$ Minimum global training loss
$R$ Number of communication rounds
@@
$\sigma_{L}^{2}$ Variance bound of local gradient estimator
$\sigma_{G}^{2}$ Dissimilarity bound between local and global gradients
$K$ Number of clients
$N$ Each global model parameter is digitalized as an $N$-bit sequence
$D$ Dimension of model $\mathbf{w}$
$T$ Total number of local training steps, $T=RE$
$L$ $L$-smoothness parameter
@@
$Q_{r}$ Maximum number of LDPC decoding iterations in the $r$-th communication round
$M$ Bound of $\max(\mathbf{w}_{r})-\min(\mathbf{w}_{r}),r=0,\cdots,R-1$
(a) Wireless FL system: Multiple mobile clients are coordinated by a remote server through noisy wireless channels, where model updates and model weights are communicated in the uplink and downlink, respectively.
(b) System operations: The server performs model aggregation, digitalization, LDPC encoding, and model broadcasting. Mobile clients perform LDPC decoding, local training, and upload the model updates. Procedures of uploading model updates are omitted.
Figure 1: Wireless FL with digital communication and its operations.
@@ II-A Wireless FL
Fig. 1(a) depicts a wireless FL system, where model training of different clients is coordinated by a remote server. The FL problem can be formulated as follows [4]:
$\min_{\mathbf{w}\in{\mathbb{R}^{D}}}f(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}f_{k}(\mathbf{w}),$ (1)
@@
where $\mathbf{w}\in{\mathbb{R}^{D}}$ denotes the DNN model with $D$ parameters, $K$ represents the number of mobile clients, $f_{k}\left(\cdot\right)$ is the local loss function of the $k$-th client, and $f\left(\cdot\right)$ is the global loss function.
Figure 2: Digitalization for each parameter in $\tilde{\mathbf{w}}_{r}$ (an example with $N=3$), where a continuous model parameter is mapped to the binary sequence of the nearest boundary.
@@
To train a shared global model, the server communicates with the clients for multiple rounds through noisy wireless channels. As shown in Fig. 1(b), in each communication round, the global model at the server (denoted as $\mathbf{w}_{r}$ for the $r$-th communication round) is broadcast to the clients. Each client then performs local training based on the received global model and transmits the model updates after local training to the server for aggregation. Specifically, in the $r$-th communication round, the global model is digitalized by an $N$-bit sequence, encoded by a channel code (e.g., the LDPC codes), and broadcast to clients. The parameter digitalization scheme is illustrated in Fig. 2, where the range of [$\min(\mathbf{w}_{r}),\max(\mathbf{w}_{r})$] is evenly divided into $12^{N}-$ 1 intervals and the width of each interval is $\Delta_{r}=\frac{\max(\mathbf{w}_{r})-\min(\mathbf{w}_{r})}{2^{N}-1}$. Each of the $2^{N}$ boundaries of these intervals are marked by a $N$-bit binary sequence. A parameter of $\mathbf{w}_{r}$ is mapped to the binary sequence of the nearest boundary. We assume $N$ is large enough so the continuous and digital representations of model parameters have negligible difference. The model parameters received/decoded by the $k$-th client at the $r$-th communication round is given as follows:
$\mathbf{w}^{k}_{r,0}=\mathbf{\tilde{w}}_{r},\ k=1,\cdots,K,\ r=0,\cdots,R-1,$ (2)
@@
where $\mathbf{\tilde{w}}_{r}$ denotes the distorted version of $\mathbf{w}_{r}$ due to channel decoding errors. To reconstruct $\tilde{\mathbf{w}}_{r}$ from the decoded bits at clients, the server also broadcasts the scalars $\min(\mathbf{w}_{r})$ and $\max(\mathbf{w}_{r})$ through an error-free downlink channel. The impact of channel decoding errors can be quantified by considering the different error bit positions in the $N$-bit representations of model parameters, which will be analyzed in Section III-A.
Each client updates the received global model $\tilde{\mathbf{w}}_{r}$ by conducting $E$ steps of local mini-batch stochastic gradient descent (SGD) with a learning rate $\eta$ as follows:
$\begin{split}\mathbf{w}^{k}_{r,e+1}=\mathbf{w}^{k}_{r,e}-\eta&{\nabla}f_{k}(\mathbf{w}^{k}_{r,e};\xi_{r,e}^{k}),\\ &e=0,\cdots,E-1,\ k=1,\cdots,K,\end{split}$ (3)
@@
where ${\nabla}f_{k}(\mathbf{w}^{k}_{r,e};\xi_{r,e}^{k})$ denotes the gradient computed from a mini-batch datasets $\xi_{r,e}^{k}$ at the $k$-th client. By completing $E$ steps of local training, the local model update of the $k$-th client can be expressed as follows:
$\begin{split}\Delta\mathbf{w}^{k}_{r}=\mathbf{w}^{k}_{r,E}-&\mathbf{w}^{k}_{r,0},\\ &r=0,\cdots,R-1,\ k=1,\cdots,K,\end{split}$ (4)
@@
$\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{K}\sum_{k=1}^{K}\Delta\mathbf{\tilde{w}}^{k}_{r},\ r=0,\cdots,R-1.$ (5)
In this work, we focus on the downlink communication rather than the uplink since our objective is to design novel channel decoding schemes to save the energy consumption of mobile clients. For the uplink, the channel decoding happens at the server, where a powerful channel decoder as well as a sufficient number of decoding iterations can always be applied to achieve satisfactory BER performance. Hence, we consider $\Delta\mathbf{\tilde{w}}_{r}^{k}=\Delta\mathbf{w}_{r}^{k}$ in the following. Define $\bar{\mathbf{w}}_{r,e}\triangleq\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{r,e}^{k}$ as the average of client models after $e$ local training steps in the $r$-th communication round. After $R$ communication rounds, the wireless FL system obtains $\bar{\mathbf{w}}_{R,E}$ as the learned model.
@@ II-B Energy Consumption Evaluation
There are two main energy-consuming operations for clients participating in wireless FL, namely model training and model communication. For the latter, it is believed that the energy is mostly consumed by the data transmission. In this subsection, we evaluate the energy consumption of different components, taking the training of AlexNet for image classification tasks as an example [20]. The analysis is based on actual measurements obtained from fabricated chips published in recent years [21, 22, 23]. For fair comparisons, all data are collected for chips fabricated in the 65 nm complementary metal-oxide semiconductor (CMOS) process.
In[21], a fabricated AI hardware accelerator was reported to achieve the energy efficiency of 13.7 mJ/epoch for 8-bit quantized AlexNet. With five epochs of local training, the processing energy consumption in each communication round is 68.5 mJ. In [22], a fifth-generation (5G) New Radio wireless transmitter and receiver were reported to achieve the energy efficiencies of 47.9 pJ/bit and 33.3 pJ/bit for transferring uncoded data, respectively, under a communication distance of around 200 meters. Therefore, the energy required to transmit and receive the 8-bit AlexNet with 60 million parameters in a single communication round is about 39 mJ. Assuming a code rate of 1/2, the communication energy consumption for coded transmission of the AlexNet model parameters in each round is 78 mJ. Besides, the LDPC decoder in [23] achieves the energy efficiency of 20.1 pJ/bit/iteration, which was measured with 10 decoding iterations on average. Accordingly, the energy consumption of the LDPC decoder for decoding the AlexNet model parameters is 96.5 mJ. The detailed comparison is presented in Table II, which indicates that the energy consumption of LDPC decoding accounts for a large proportion of the overall energy consumption at clients in wireless FL.
TABLE II: Energy Consumption of Main Components at a Mobile Unit in Wireless FL for Training an 8-bit AlexNet
@@
RF Transceiver 81.2 pJ/bit[22] 78 mJ
LDPC Decoder 20.1 pJ/bit/iter.[23] 96.5 mJ